OpenStudy (anonymous):

Use the change of base formula to evaluate each to three decimal places. log(8)105

OpenStudy (anonymous):

Do you know the change of base formula and where it comes from? if we rewrite this as an exponential \[\log_{8}105\rightarrow 8^{x}=105\]. Now take the logarithm of both sides in base 10 or e: \[\ln \left( 8^{x} \right)=\ln \left( 105 \right)\rightarrow xln\left( 8 \right)=\ln \left( 105 \right) \rightarrow x=\frac{ \ln \left( 105 \right) }{ \ln\left( 8 \right) }\] You can use a calculator to compute these.

OpenStudy (anonymous):

so then it would be 2.238?

OpenStudy (anonymous):

\[x=\ln _{8}105\]

OpenStudy (anonymous):

oops, I mean \[x=\log_{8} 105\] This is 2.238... you did great!